Write down centre of mass semi circular ring.
1
Expert's answer
2020-09-15T10:05:17-0400

solution


consider a semicircular ring of radius R which is uniform.a uniform semicircular ring is symmetric about Y-axis so its its x coordinate of center of mass will be zero .


diagram for this question can be drawn as




y coordinate of the semicircular ring can be written as


"y=\frac{2R}{\pi}"


therefore center of mass is at "(0,\frac{2R}{\pi})"